OpenStudy (anonymous):

Solve using the Quadratic Formula. x² - 6x - 7 = 0

OpenStudy (btaylor):

\[x=\frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \rightarrow x=\frac{6 \pm \sqrt{36+28}}{2} \rightarrow x = \frac{6 \pm \sqrt{64}}{2}\] so you get (6+8)/2 = 7 and (6-8)/2 = -1 so x = 7 and x = -1

OpenStudy (anonymous):

a = 1 b= -6 c = -7 Use: \(D = b^2 - 4ac\) D = 36 + 28 = 64 \[\sqrt{D} = 8\] Now use: \[x = \frac{6 \pm 8}{2}\] x = 7 and x = -1

OpenStudy (anonymous):

Hey, Jeletta. I can help you. Do you have any specific questions about how to do this? The basic steps are: 1) Write the equation set equal to 0. (If it's not already) 2) Figure out what A, B, and C are. 3) Plug into the formula and simplify it.